• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. LOPEZ

    Log entry time 09:30:12 on December 01, 2007

    Entry number 217469

    keyword=BPM POSITION FOR LAST RUNS

    BPM position for He4 runs 3117 (LEFT) and 21337 (RIGHT) was about (-1,1). It should be (-1,0). We found it after having changed the momentum of both spectrometers. We called MCC to fix it.

    After talking with the RC, we decided to go on with the new momentum setting instead of changing the momentum back.

    At some point it should be measured again...